# Marc Martinez-Gost12, Ana Pérez-Neira123 1Centre Tecnològic de Telecomunicacions de Catalunya, Spain 2Dept. of Signal Theory and Communications, Universitat Politècnica de Catalunya, Spain 3ICREA Acadèmia, Spain

### v1

@@ Abstract
This paper proposes a novel split learning architecture designed to exploit the cyclical movement of Low Earth Orbit (LEO) satellites in non-terrestrial networks (NTNs). Although existing research focuses on offloading tasks to the NTN infrastructure, these approaches overlook the dynamic movement patterns of LEO satellites that can be used to efficiently distribute the learning task. In this work, we analyze how LEO satellites, from the perspective of ground terminals, can participate in a time-window-based model training. By splitting the model between a LEO and a ground terminal, the computational burden on the satellite segment is reduced, while each LEO satellite offloads the partially trained model to the next satellite in the constellation. This cyclical training process allows larger and more energy-intensive models to be deployed and trained across multiple LEO satellites, despite their limited energy resources. We formulate an optimization problem that manages radio and processing resources, ensuring the entire data is processed during each satellite pass while minimizing the energy consumption. Our results demonstrate that this approach offers a more scalable and energy-efficient way to train complex models, enhancing the capabilities of LEO satellite constellations in the context of Artificial Intelligence-driven applications.
@@ I Introduction
The last two decades have seen an unprecedented growth towards the deployment of satellite mobile edge computing (SMEC) over low Earth orbit (LEO) satellite constellations [1]. SMEC extends traditional MEC concepts to satellite networks and comprises exploiting the computational resources (e.g., servers, GPUs, or specialized processing units) on satellites to perform tasks like data processing and caching. This trend can be attributed to several key technological and economic advancements. Specifically, modern satellites are now equipped with advanced capabilities, integrating cutting-edge communication, computing, and sensing technologies. These enhancements allow LEO satellites to not only provide traditional services, such as global communication and Earth observation, but also enable them to support more sophisticated applications, including distributed computing and Artificial Intelligence (AI) model inference closer to the data source or end users.
@@ II Edge Learning over LEO networks
Federated learning (FL) [7, 8, 9] has emerged as a promising technique for distributed AI model training in Native SMEC, particularly due to its ability to handle the decentralized nature of satellite data. In FL, instead of sending raw data to a central server, satellites locally train models and share only updates, preserving data privacy and reducing the need for large-scale data transmissions. However, FL presents several challenges in the satellite context. First, communication overhead for aggregating updates can consume significant bandwidth and energy. As satellite constellations grow, the communication burden increases, potentially leading to inefficiencies. Second, while FL distributes training across multiple entities, it may still require large models to achieve high performance for complex tasks. Such large models are not energy-preserving, as training and updating them on satellites with constrained power resources can be prohibitively expensive. Thus, while FL is well-suited for distributed AI in satellite networks, addressing communication, computational, and energy constraints is critical to its success in this context.
@@ III System Model
We consider a set of LEO satellites capturing data with its sensors (e.g., optical images, hyperspectral data, SAR, etc.) with the goal of training an algorithm (e.g, a neural network). Due to the overall amounts of data and the limited communication bandwidth, sending the raw data towards the ground for processing is not feasible. On the other hand, running the algorithm at the satellite premises is complex because it is energy-intensive and the data is distributed across the LEO constellation. Following the SL paradigm, the algorithm is split into two parts; the first piece is deployed at the LEO and the latter at the ground terminal. During the satellite pass of a specific LEO, the satellite and the ground terminal collaboratively execute the algorithm, leveraging their communication and computational capabilities. Once the satellite pass concludes, the satellite hands off the split to the following satellite in the orbital ring. This succeeding satellite will continue the training process using its local data. Figure 1 illustrates this joint communication and processing procedure. The proposed scenario can be formulated in the opposite direction, this is, a ground terminal with stringent energy constraints exploiting the communication capabilities with a set of LEO satellites.
Figure 1: Procedure of the proposed SL architecture. After step (8), the whole process is repeated between $n+1$-th satellite and the ground terminal.
@@ III-A Network architecture
For the sake of simplicity we assume the ground terminal communicates with a single LEO satellite during the satellite pass. All satellites belong to the same orbital ring, which is comprised of $N$ evenly distributed satellites deployed at altitude $h$. We will determine the duration of each satellite pass over the ground terminal, as well as the distances for the ground-to-satellite link (used for split learning communication) and the inter-satellite links (used to exchange the satellite’s portion of the model). These metrics are essential for understanding the communication and timing constraints of the proposed system.
@@ III-C Communication
where $D_{\text{ISL}}$ corresponds to the size of the split in bits and $R_{\text{ISL}}$ is the ISL data rate. The associated propagation time is $\bar{d}/{c}$ and the spent energy is $E_{\text{ISL}}=p_{\text{tx}}^{\text{ISL}}~{}T_{\text{ISL}}$, where $p_{\text{tx}}^{\text{ISL}}$ is the power used in to transmit in the ISL.
IV Energy Minimization in Split Learning
@@
$E_{\text{total}}=$ $~{}E_{\text{proc}}(W_{1},f_{p}^{\text{GS}})+E_{\text{comm}}(p_{\text{tx}}^{\text{GS}})+E_{\text{proc}}(W_{2},f_{p}^{\text{LEO}})$
$+E_{\text{comm}}(p_{\text{tx}}^{\text{LEO}})+E_{\text{ISL}}(p_{\text{tx}}^{\text{ISL}}),$ (11)
where $f_{p}^{\text{GS}}$ and $f_{p}^{\text{LEO}}$ are the frequency of the processor at the ground terminal and at the satellite, respectively; $p_{\text{tx}}^{\text{GS}}$ and $p_{\text{tx}}^{\text{LEO}}$ are the transmission power of the ground terminal and the satellite, respectively.
@@
$T_{\text{total}}=$ $~{}T_{\text{proc}}(W_{1},f_{p}^{\text{GS}})+T_{\text{comm}}(p_{\text{tx}}^{\text{GS}})+2~{}T_{\text{prop}}$
$+T_{\text{proc}}(W_{2},f_{p}^{\text{LEO}})+T_{\text{comm}}(p_{\text{tx}}^{\text{LEO}})+T_{\text{ISL}},$ (12)
@@ V-A Image compression
For this implementation, we use an autoencoder to compress color images from $224\times 224$ pixels down to a $7\times 7$ latent space. The encoder is allocated at the satellite premises, while the decoder is located at the ground terminal. Assuming 32 bits per datum, the LEO satellite transmits $D_{\text{tx}}=4.7$ kbits per image (with the same size assumed for the gradients in the uplink). Additionally, the encoder weighs $D_{\text{ISL}}=168.8$ kbits. We use the fvcore library [17] to estimate the load of the algorithm. The encoder requires $W_{1}\approx 302$ GFLOPS and the decoder requires $W_{2}\approx 39$ MFLOPS.
@@
Figure 3: Energy consumption: (Top) Autoencoder with SL and direct download; (Bottom) ResNet-18 at three different splitting points.
@@ V-B Image classification
Image classification of satellite imagery is crucial for applications where quick and accurate analysis of large-scale images is essential for monitoring vast areas. In such scenarios, SL provides a key advantage by preserving the privacy of the data, as raw images are never transmitted. We use a ResNet-18 architecture, which is a lightweight and widely-used neural network for classification tasks [18]. Despite being efficient, the overall computational cost of the ResNet model is higher compared to the autoencoder used in the previous task, due to the complexity of the classification task and the deeper layers involved. The work of the algorithm and data sizes for three different splitting points are listed in Table II.
Figure 3 (Bottom) illustrates how varying the splitting point within the ResNet-18 architecture affects the overall energy consumption of the system. We explore three different partitioning points, each one leaving more work to the satellite. For this task, splitting the model to process more at the satellite premises proves to be advantageous, as the output generated at the satellite is smaller, reducing the amount of data transmitted and therefore minimizing energy spent on communication and ground processing.

### v2

@@ Abstract
This paper proposes a split learning (SL) framework tailored for Low Earth Orbit (LEO) satellite constellations, leveraging their cyclical movement to improve energy efficiency. Although existing research focuses on offloading tasks to the non-terrestrial network (NTN) infrastructure, these approaches overlook the dynamic movement patterns of LEO satellites that can be used to efficiently distribute the learning task. In this work, we analyze how LEO satellites, from the perspective of ground terminals, can participate in a time-window-based model training. By splitting the model between a LEO and a ground terminal, the computational burden on the satellite segment is reduced, while each LEO satellite offloads the partially trained model to the next satellite in the constellation. This cyclical training process allows larger and more energy-intensive models to be deployed and trained across multiple LEO satellites, despite their limited energy resources. We formulate an optimization problem that manages radio and processing resources, ensuring the entire data is processed during each satellite pass while minimizing the energy consumption. Results demonstrate that the proposed architecture optimizes both communication and processing resources, achieving up to 97% energy savings compared to direct raw data transmission. This approach offers a more scalable and energy-efficient way to train complex models, enhancing the capabilities of LEO satellite constellations in Artificial Intelligence-driven applications.
@@ I Introduction
The last two decades have seen an unprecedented growth towards the deployment of satellite mobile edge computing (SMEC) over low Earth orbit (LEO) satellite constellations [1]. SMEC extends traditional MEC concepts to satellite networks and comprises exploiting the computational resources (e.g., servers, GPUs, or specialized processing units) on satellites to perform tasks like data processing and caching. This trend can be attributed to several key technological and economic advancements. Specifically, modern satellites are now equipped with advanced capabilities, integrating cutting-edge communication, computing, and sensing technologies. These enhancements allow LEO satellites not only to provide traditional services, such as global communication and Earth observation, but also enable them to support more sophisticated applications, including distributed computing and Artificial Intelligence (AI) model inference closer to the data source or end users.
@@ II Edge Learning over LEO networks
Federated learning (FL) [7, 8, 9] has emerged as a promising technique for distributed AI model training in Native SMEC, particularly due to its ability to handle the decentralized nature of satellite data. In FL, instead of sending raw data to a central server, satellites train models locally and share only updates, preserving data privacy and reducing the need for large-scale data transmissions. However, FL presents several challenges in the satellite context. First, communication overhead for aggregating updates can consume significant bandwidth and energy. As satellite constellations grow, the communication burden increases, potentially leading to inefficiencies. Second, while FL distributes training across multiple entities, it may still require large models to achieve high performance for complex tasks. Such large models are not energy-preserving, as training and updating them on satellites with constrained power resources can be prohibitively expensive. Thus, while FL is well-suited for distributed AI in satellite networks, addressing communication, computational, and energy constraints is critical to its success in this context.
@@ III System Model
We consider a set of LEO satellites capturing data with its sensors (e.g., optical images, hyperspectral data, SAR, etc.) with the goal of training an algorithm (e.g, a neural network). Due to the overall amounts of data and the limited communication bandwidth, sending the raw data towards the ground for processing is not feasible. On the other hand, running the algorithm at the satellite premises is complex because it is energy-intensive and the data is distributed across the LEO constellation. Following the SL paradigm, the algorithm is split into two parts: the first piece is deployed at the LEO and the latter at the ground terminal. During the satellite pass of a specific LEO, the satellite and the ground terminal collaboratively execute the algorithm, leveraging their communication and computational capabilities. Once the satellite pass concludes, the satellite hands off the split to the following satellite in the orbital ring. This succeeding satellite will continue the training process using its local data. Figure 1 illustrates this joint communication and processing procedure. The proposed scenario can be formulated in the opposite direction, this is, a ground terminal with stringent energy constraints exploiting the communication capabilities with a set of LEO satellites.
Figure 1: Procedure of the proposed SL architecture. After step (8), the whole process is repeated between $(n+1)$-th satellite and the ground terminal.
@@ III-A Network architecture
For the sake of simplicity we assume the ground terminal communicates with a single LEO satellite during the satellite pass. All satellites belong to the same orbital ring, which is comprised of $N$ evenly distributed satellites deployed at altitude $h$. We will determine the duration of each satellite pass over the ground terminal, as well as the distances for the ground-to-satellite link (used for SL communication) and the inter-satellite links (used to exchange the satellite’s portion of the model). These metrics are essential for understanding the communication and timing constraints of the proposed system.
@@ III-C Communication
where $D_{\text{ISL}}$ corresponds to the size of the split in bits and $R_{\text{ISL}}$ is the ISL data rate. The associated propagation time is $d_{\text{ISL}}/{c}$ and the spent energy is $E_{\text{ISL}}=p_{\text{tx}}^{\text{ISL}}~{}T_{\text{ISL}}$, where $d_{\text{ISL}}$ is the ISL distance and $p_{\text{tx}}^{\text{ISL}}$ is the power used in to transmit in the ISL.
IV Energy Minimization in SL
@@
$E_{\text{total}}=$ $~{}E_{\text{proc}}(W_{1},f_{p}^{\text{LEO}})+E_{\text{comm}}(p_{\text{tx}}^{\text{LEO}})+E_{\text{proc}}(W_{2},f_{p}^{\text{GS}})$
$+E_{\text{comm}}(p_{\text{tx}}^{\text{GS}})+E_{\text{ISL}}(p_{\text{tx}}^{\text{ISL}}),$ (11)
where $f_{p}^{\text{LEO}}$ and $f_{p}^{\text{GS}}$ are the frequency of the processor at the satellite and ground terminal, respectively; $p_{\text{tx}}^{\text{LEO}}$ and $p_{\text{tx}}^{\text{GS}}$ are the corresponding transmission powers.
@@
$T_{\text{total}}=$ $~{}T_{\text{proc}}(W_{1},f_{p}^{\text{LEO}})+T_{\text{comm}}(p_{\text{tx}}^{\text{LEO}})+2~{}T_{\text{prop}}$
$+T_{\text{proc}}(W_{2},f_{p}^{\text{GS}})+T_{\text{comm}}(p_{\text{tx}}^{\text{GS}})+T_{\text{ISL}},$ (12)
@@ V-A Image compression
For this implementation, we use an autoencoder to compress color images from $224\times 224$ pixels down to a $7\times 7$ latent space. The encoder is allocated at the satellite premises, while the decoder is located at the ground terminal. Assuming 32 bits per datum, the LEO satellite transmits $D_{\text{tx}}=4.7$ kbits per image (with the same size assumed for the gradients in the uplink). Additionally, the encoder weighs $D_{\text{ISL}}=168.8$ kbits. We use the fvcore library [17] to estimate the load of the algorithm. The encoder and decoder require $W_{1}\approx 302$ GFLOPS and $W_{2}\approx 39$ MFLOPS, respectively.
@@
Figure 3: Energy consumption: (Top) Autoencoder with SL and direct download; (Bottom) ResNet-18 at three different splitting points $\ell$.
@@ V-B Image classification
Image classification of satellite imagery is crucial for applications where quick and accurate analysis of large-scale images is essential for monitoring vast areas. In such scenarios, SL provides a key advantage by preserving the privacy of the data, as raw images are never transmitted. We use a ResNet-18 architecture, which is a lightweight and widely-used neural network for classification tasks [18]. Despite being efficient, the overall computational cost of the ResNet model is higher compared to the autoencoder used in the previous task, due to the complexity of the classification task and the deeper layers involved. The work of the algorithm and data sizes for three different splitting points $\ell$ are listed in Table II.
Figure 3 (Bottom) illustrates how varying the splitting point within the ResNet-18 architecture affects the overall energy consumption of the system. For this task, splitting the model to process more at the satellite premises (i.e., $\ell_{3}$) proves to be advantageous, as the satellite’s output is smaller, reducing the communication cost and, thus, minimizing the energy spent on communication and ground processing. These results show that the optimal splitting point depends on the learning model architecture, demanding tailored solutions for each use case.